# Efficient Learning with Sine-Activated Low-Rank Matrices

### v4

@@ Abstract
Low-rank decomposition has emerged as a vital tool for enhancing parameter efficiency in neural network architectures, gaining traction across diverse applications in machine learning. These techniques significantly lower the number of parameters, striking a balance between compactness and performance. However, a common challenge has been the compromise between parameter efficiency and the accuracy of the model, where reduced parameters often lead to diminished accuracy compared to their full-rank counterparts. In this work, we propose a novel theoretical framework that integrates a sinusoidal function within the low-rank decomposition process. This approach not only preserves the benefits of the parameter efficiency characteristic of low-rank methods but also increases the decomposition’s rank, thereby enhancing model performance. Our method proves to be a plug in enhancement for existing low-rank models, as evidenced by its successful application in Vision Transformers (ViT), Large Language Models (LLMs), Neural Radiance Fields (NeRF) and 3D shape modelling.
@@ 1 Introduction
We direct the reader’s attention to figure 1, which showcases our method across various machine learning tasks. Our comparisons with standard low-rank methods consistently demonstrate superior performance across these diverse tasks.
@@
Figure 1: Applying a drop-in sine-activation, sine low-rank (ours), component increases the rank of low-rank matrix methods, leading to improved parameter efficient performance on a variety of tasks including: a) NeRF, b) 3D Occupancy, c) ViT image classification, and d) Fine-tuning Large Language Models (LoRA).
Our approach’s advantages are corroborated across a range of machine learning applications, including variations of LoRA methods for ViT (He et al., 2022b), LLMs (Hu et al., 2022), NeRF for novel view synthesis (Mildenhall et al., 2020), and 3D shape modeling via Binary Occupancy Fields (Mescheder et al., 2019). Across the board, our approach not only matches the parameter savings offered by low-rank methods but also results in an improvement in accuracy, showing its broad applicability and superior performance among diverse machine learning tasks. The main contributions of our paper are:
@@
    A theoretical framework demonstrating that applying sinusoidal non-linearities to low-rank matrices can effectively increase their rank without introducing additional parameters. This method leverages the inherent properties of sinusoidal functions to augment the expressiveness of low-rank matrices, allowing them to capture more complex patterns while maintaining parameter efficiency.
@@ 2.1 Low-rank decomposition:
Low-rank decomposition stands as a crucial method across disciplines like information theory, optimization, and machine learning, providing a strategic approach to reduce memory costs (Strang, 2019; Xinwei et al., 2023). Notably, (Candès et al., 2011) uncovered that matrices can precisely separate low-rank and sparse components through convex programming, linking to matrix completion and recovery. Expanding its application, (Yu et al., 2017) devised a low-rank learning framework for Convolutional Neural Networks, enhancing compression while maintaining accuracy. (Sharma et al., 2023) further found that performance improvements in Large Language Models could be achieved by eliminating higher-order weight matrix components without extra parameters or data. In the realm of neural radiance fields, (Tang et al., 2022) introduced a rank-residual learning strategy for optimal low-rank approximations, facilitating model size adjustments. Additional contributions include (Shi & Guillemot, 2023) with rank-constrained distillation, (Chen et al., 2022) applying vector-matrix decomposition, and (Schwarz et al., 2023) using soft-gated low-rank decompositions for compression. More recently, (Yuan & Zhao, 2023) implemented a vector-matrix decomposition strategy that allows for test-time compression adjustments.
2.2 Parameter efficient learning:
Parameter efficient learning is an important research area in deep learning, merging various techniques to enhance model adaptability with minimal resource demands (Menghani, 2023). Techniques like parameter-efficient fine-tuning (PEFT) allow pretrained models to adjust to new tasks efficiently, addressing the challenges of fine-tuning large models due to high hardware and storage costs. Among these, Visual Prompt Tuning (VPT) stands out for its minimal parameter alteration—less than 1%—in the input space, effectively refining large Transformer models while keeping the core architecture unchanged (Jia et al., 2022). Similarly, BitFit offers a sparse-finetuning approach, tweaking only the model’s bias terms for cost-effective adaptations (Zaken et al., 2022). Moreover, LoRA introduces a low-rank adaptation that maintains model quality without additional inference latency or altering input sequence lengths, by embedding trainable rank matrices within the Transformer layers (Hu et al., 2022). Recent studies also combine LoRA with other efficiency strategies like quantization, pruning, and random projections for further model compression (Dettmers et al., 2024; Li et al., 2024; Zhang et al., 2024; Kopiczko et al., 2024)
@@ 3 Methodology
We introduce our technique that we term a sine activated low-rank approximation. The main purpose of this technique is to increase the rank of an initial low-rank matrix without adding any extra parameters.
@@ 3.1.1 Feed-Forward Layer
where $\mathbf{W}\,{\in}\,\mathbb{R}^{m{\times}n}$ is a dense weight matrix, $\mathbf{b}\,{\in}\,\mathbb{R}^{m{\times}1}$ is the bias of the layer, and $\mathbf{x}$ is the input from the previous layer. The output $\mathbf{y}$ is then often activated by a non-linearity $\sigma$ producing $\sigma(\mathbf{y})$. The weight matrix $\mathbf{W}$ and bias $\mathbf{b}$ are trainable parameters of the layer. In contemporary deep learning models, the feed-forward layers’ weight matrices, $\mathbf{W}$, are often large and dense yielding a high rank matrix. While the high-rank property of the weight matrix helps in representing complex signals, it significantly adds to the overall parameter count within the network yielding the need for a trade-off between the rank of the weight matrix and overall architecture capacity.
@@ 3.1.2 Low-rank decomposition
This is the most common way to reduce the parameter count in the feed-forward layer. During the training process, this method performs optimization on $\mathbf{U}$ and $\mathbf{V}$ alternatively. Low-rank multiplication then reduces the learnable parameter count and memory footprint from $O(mn)$ to $O(k\cdot(m+n))$. Although $\mathbf{U}\mathbf{V}^{T}$ has the same matrix shape as the full-rank matrix $\mathbf{W}$, the rank of $\mathbf{U}\mathbf{V}^{T}$ is constrained and $rank(\mathbf{U}\mathbf{V}^{T})\leq r$. Thus while we have significantly decreased the number of trainable weights in such a layer, we have paid the price by obtaining a matrix of much smaller rank. In the next section, we address this trade-off by developing a technique that can raise back the rank of a low-rank decomposition while keeping its low parameter count.
@@ 3.2.1 Non-linear low-rank decomposition
where $\phi(\cdot)$ is the non-linearity function, $\omega$ is a non-learnable frequency parameter, and $g$ is a non-learnable parameter to control the gain of the transformation. Our theoretical work will show that $\phi(\omega x)=sin(\omega x)$ is a very good choice for the non-linearity $\phi$.
@@ 3.2.2 Main result
The following theorem relates the rank of $\sin(\omega\cdot\mathbf{A})$ to the frequency parameters $\omega$ and the quantity $A_{\min}^{0}$.
@@
We mention that theorem 1 also holds for the case where we initialize $\mathbf{U}$ and $\mathbf{V}$ by a normal distribution of variance $N$.
Weight matrices within feed-forward layers are typically initialized using a distribution that is contingent upon the layer’s neuron count. When considering low-rank decompositions characterized by matrices $\mathbf{U}\in\mathbb{R}^{m\times k}$ and $\mathbf{V}^{T}\in\mathbb{R}^{k\times n}$, where $k\ll\min\{m,n\}$, the variance of this initialization distribution is influenced by $m$ and $n$. These dimensions are significantly larger than $k$, ensuring that the condition specified in theorem 1 — that $N>k$ — is always met, making this theorem especially relevant for low-rank decompositions in feed-forward layers. For example the most common initialization schemes such as Kaiming (He et al., 2015b) and Xavier (Glorot & Bengio, 2010) satisfy the requirements of our theorem.
@@
Figure 2: These figures display weight magnitudes for matrices with dimension $128\times 128$. The first figure showcases a heatmap of a full-rank matrix initialized by Kaiming uniform, highlighting linear independence among rows. The second shows a low-rank matrix $\mathbf{W_{\text{lr}}}\,{=}\,\mathbf{UV}^{T}\,{\in}\,\mathbb{R}^{128{\times}128}$, with $\mathbf{U},\mathbf{V}\,{\in}\,\mathbb{R}^{128{\times}1}$ initialized by Kaiming uniform illustrating minimal linear independence. The final pair of figures reveal how applying a sine function element-wise, $\sin(\omega\cdot\mathbf{UV}^{T})$, with varying $\omega$, affects linear independence in low-rank matrices; specifically, $\omega=100$ and $\omega=2000$ progressively increase linear independence.
In figure 2 we give a visualization of our method in action. We consider a full-rank matrix, a low-rank matrix, and two sine activated low-rank matrices with different frequencies. By visualizing the weight magnitudes in each matrix via a heat map, we can clearly see how the sine activated low-rank matrix increases rank and furthermore how increasing the frequency of the sine function increases the rank in accord with theorem 1.
@@ 4.1 Large Language Model
Setting. In the Transformer architecture, there are four weight matrices in the self-attention module ($\mathbf{W_{\text{q}}},\mathbf{W_{\text{k}}},\mathbf{W_{\text{v}}},\mathbf{W_{\text{o}}}$) and two in the MLP module($\mathbf{W_{\text{up}}},\mathbf{W_{\text{down}}}$). To evaluate RoBERTA V3, we follow up the LoRA architecture and implement low-rank adaptation only on $\mathbf{W_{\text{q}}}$ and $\mathbf{W_{\text{v}}}$. We study the performance of LoRA and sine LoRA in terms of different rank $k={1,2,4,8}$. To evaluate LLaMA 3-8B, we implement low-rank adaptations on five modules ($\mathbf{W_{\text{q}}},\mathbf{W_{\text{k}}},\mathbf{W_{\text{v}}},\mathbf{W_{\text{up}}},\mathbf{W_{\text{down}}}$) with different rank $k={4,8,16,32}$. For further details on implementation and computational cost please see Section A.2.1.
@@
Analysis. Within the LoRA framework, featuring a low-rank multiplication component $\Delta\mathbf{W}=\mathbf{UV^{T}}$, we enhance this low-rank component with a sine function and assess the efficacy of our method. This adaptation amplifies the update significance due to the ‘intrinsic rank’ increase, facilitated by the sine activation. Consequently, our approach attains superior performance at reduced rank levels $k$, compared to LoRA, effectively decreasing the count of learnable parameters.
@@
For comparisons of our method on the recent DORA paper Liu et al. (2024) please see Section A.2.1.
@@ 4.2 Pretraining Vision Transformers
Figure 4: Low Rank ViT classification performance. Use of the sine activation improves performance of the low-rank models, and even enables improvement relative to the Full Rank model.
@@
Analysis. Large models, such as ViT-Base, with an excessively large number of parameters, are prone to overfitting, where they perform well on training data but poorly on unseen data, especially when trained on relatively ‘small’ datasets like ImageNet-1k (Zheng et al., 2024; Xu et al., 2024). Low-rank learning techniques can help in designing models that generalize better to new data by encouraging the model to learn more compact and generalizable representations to reduce overfitting. Additionally, while ViT architectures often underperform on smaller datasets, this method introduces a novel approach for efficiently training ViT models using small data collections. For frequency ablations in the rank 1 case, see Section A.2.2.
@@ 4.3 NeRF
Figure 5: (a) Using a non-transformed Low-Rank model leads to a complete loss of signal at extreme (rank $k=1$). In contrast, applying a sine activation function is able to reconstruct details even at 1.3% of the Full-Rank parameters. (b) The Sine Low-Rank NeRF models show significant improvements across the rate-distortion curve relative to the Low-Rank models.
Analysis. NeRF models fit entire 3D scenes, and a high training PSNR leads to a high testing PSNR (Mildenhall et al., 2020). Employing structured weight matrices could result in a drop in performance due to the inherent constraints imposed by their structural design. Increasing the matrices’ rank enhances their memorization abilities significantly, especially when using a very low $k$. Starting from a low frequency, there is a rapid and consistent increase in PSNR. Consequently, as we elevate the rank level $k$, our results gradually align with the baseline NeRFs, which serve as the upper bound. For ablations on the choice of frequency in the case of rank 1 and rank 5 see Section A.2.4.
@@
For this experiment, we evaluate performance on binary occupancy field reconstruction, which involves determining whether a given coordinate is occupied (Mescheder et al., 2019). Following (Saragadam et al., 2023), we sampled over a 512 $\times$ 512 $\times$ 512 grid with each voxel within the volume assigned a 1, and voxels outside the volume assigned a 0. We use the Thai Statue, Dragon and Lucy instance from the Stanford Scanning Repository.***Available at https://graphics.stanford.edu/data/3Dscanrep/ We evaluate intersection over union (IoU) for the occupancy volumes. We used a coordinate-based MLP that includes two hidden layers, each with a width of 256 neurons, and employed the Gaussian activation function. The full-rank model achieves an accuracy of 97 (IoU). Figure 6 shows the 3D mesh representation of the Thai Statue, visualized using the low-rank method and the sine low-rank method for $k=1,2,5$. Applying the sine function to the low-rank matrix resulted in a significant enhancement and more precise shape delineation. Results on Dragon and Lucy are given in Table 13 in Section A.2.5. A frequency ablation in the rank 1 case is given in Section A.2.5.

### v5

@@ Abstract
Low-rank decomposition has emerged as a vital tool for enhancing parameter efficiency in neural network architectures, gaining traction across diverse applications in machine learning. These techniques significantly lower the number of parameters, striking a balance between compactness and performance. However, a common challenge has been the compromise between parameter efficiency and the accuracy of the model, where reduced parameters often lead to diminished accuracy compared to their full-rank counterparts. In this work, we propose a novel theoretical framework that integrates a sinusoidal function within the low-rank decomposition. This approach not only preserves the benefits of the parameter efficiency of low-rank methods but also increases the decomposition’s rank, thereby enhancing model performance. Our method proves to be a plug-in enhancement for existing low-rank methods, as evidenced by its successful application in Vision Transformers (ViT), Large Language Models (LLMs), Neural Radiance Fields (NeRF) and 3D shape modelling. The code is publicly available at https://samy-ji.github.io/sine_activated_PEL/.
@@ 1 Introduction
We direct the reader’s attention to figure 1, which showcases our method across various machine learning applications. Our comparisons with standard low-rank methods consistently demonstrate superior performance across these diverse tasks.
@@
Figure 1: Applying a drop-in sine-activation increases the rank of low-rank matrix methods, leading to improved parameter efficiency and performance on a variety of tasks including: a) NeRF, b) 3D Occupancy, c) ViT image classification, and d) Fine-tuning Large Language Models (LoRA).
Our approach’s advantages are corroborated across a range of machine learning applications, including low rank methods for ViT (He et al., 2022b), LLMs (Hu et al., 2022; Liu et al., 2024), NeRF for novel view synthesis (Mildenhall et al., 2020), and 3D shape modeling via Binary Occupancy Fields (Mescheder et al., 2019). Across the board, our approach not only matches the parameter savings offered by low-rank methods but also results in an improvement in accuracy, showing its broad applicability and superior performance among diverse machine learning tasks. The main contributions of our paper are:
@@
    A theoretical framework demonstrating that applying sinusoidal non-linearities to low-rank matrices can effectively increase their rank without introducing additional parameters.
@@ 2.1 Low-rank decomposition:
Low-rank decomposition stands as a crucial method across disciplines such as information theory, optimization, and machine learning, providing an approach to reduce memory costs (Strang, 2019; Xinwei et al., 2023). Notably, (Candès et al., 2011) uncovered that matrices can precisely separate low-rank and sparse components through convex programming, linking to matrix completion and recovery. Expanding its application, (Yu et al., 2017) devised a low-rank learning framework for convolutional neural networks, enhancing compression while maintaining accuracy. (Sharma et al., 2023) found that performance improvements in large language models could be achieved by eliminating higher-order weight matrix components without extra parameters or data. In neural radiance fields, (Tang et al., 2022) introduced a rank-residual learning strategy for optimal low-rank approximations, facilitating model size adjustments. Additional contributions include (Shi & Guillemot, 2023) with rank-constrained distillation, (Chen et al., 2022) applying vector-matrix decomposition, and (Schwarz et al., 2023) using soft-gated low-rank decompositions for compression.
2.2 Parameter-efficient learning:
Parameter-efficient learning is an important research area in deep learning, merging various techniques to enhance model adaptability with minimal resource demands (Menghani, 2023). Techniques like parameter-efficient fine-tuning (PEFT) allow pretrained models to adjust to new tasks efficiently, addressing the challenges of fine-tuning large models due to high hardware and storage costs. Among these, Visual Prompt Tuning (VPT) stands out for its minimal parameter alteration—less than 1%—in the input space, effectively refining large Transformer models while keeping the core architecture unchanged (Jia et al., 2022). Similarly, BitFit offers a sparse-finetuning approach, tweaking only the model’s bias terms for cost-effective adaptations (Zaken et al., 2022). Moreover, LoRA introduces a low-rank adaptation that maintains model quality without additional inference latency or altering input sequence lengths, by embedding trainable rank matrices within the Transformer layers (Hu et al., 2022). Recent studies also combine LoRA with other efficiency strategies like quantization, pruning, and random projections for further model compression (Dettmers et al., 2024; Li et al., 2024; Zhang et al., 2024; Kopiczko et al., 2024; Liu et al., 2024).
@@ 3 Methodology
We introduce our main technique that we term a sine-activated low-rank approximation. The purpose of this technique is to increase the rank of a low-rank matrix without adding any extra parameters.
@@ 3.1.1 Feed-Forward Layer
where $\mathbf{W}\,{\in}\,\mathbb{R}^{m{\times}n}$ is a dense weight matrix, $\mathbf{b}\,{\in}\,\mathbb{R}^{m{\times}1}$ is the bias of the layer, and $\mathbf{x}$ is the input from the previous layer. The output $\mathbf{y}$ is often activated by a non-linearity $\sigma$ producing $\sigma(\mathbf{y})$. The weight matrix $\mathbf{W}$ and bias $\mathbf{b}$ are trainable parameters of the layer. In contemporary deep learning models, the feed-forward layers’ weight matrices, $\mathbf{W}$, are often large and dense yielding a high rank matrix. While the high-rank property of the weight matrix helps in representing complex signals, it significantly adds to the overall parameter count within the network yielding the need for a trade-off between the rank of the weight matrix and overall architecture capacity.
@@ 3.1.2 Low-rank decomposition
This is the most common way to reduce the parameter count in a feed-forward layer. During the training process, this method performs optimization on $\mathbf{U}$ and $\mathbf{V}$ alternatively. Low-rank multiplication then reduces the learnable parameter count and memory footprint from $O(mn)$ to $O(k\cdot(m+n))$. Although $\mathbf{U}\mathbf{V}^{T}$ has the same matrix shape as the full-rank matrix $\mathbf{W}$, the rank of $\mathbf{U}\mathbf{V}^{T}$ is constrained and $rank(\mathbf{U}\mathbf{V}^{T})\leq k$. Thus while we have significantly decreased the number of trainable weights in such a layer, we have paid the price by obtaining a matrix of much smaller rank. In the next section, we address this trade-off by developing a technique that can raise back the rank of a low-rank decomposition while keeping its low parameter count.
@@ 3.2.1 Non-linear low-rank decomposition
where $\phi(\cdot)$ is the non-linearity function, $\omega$ is a non-learnable frequency parameter, and $g$ is a non-learnable parameter to control the gain of the transformation. Our theoretical work will show that $\phi(\omega x)=sin(\omega x)$ is an optimal choice to increase the rank of $\mathbf{U}\mathbf{V}^{T}$.
@@ 3.2.2 Main result
The following theorem relates the rank of $\sin(\omega\cdot\mathbf{A})$ to the frequency $\omega$ and the quantity $A_{\min}^{0}$.
@@
Theorem 1 also holds for the case we initialize $\mathbf{U}$ and $\mathbf{V}$ by a normal distribution of variance $N$.
Weight matrices within feed-forward layers are typically initialized using a distribution that is contingent upon the layer’s neuron count. When considering low-rank decompositions characterized by matrices $\mathbf{U}\in\mathbb{R}^{m\times k}$ and $\mathbf{V}^{T}\in\mathbb{R}^{k\times n}$, where $k\ll\min\{m,n\}$, the variance of this initialization distribution is influenced by $m$ and $n$. These dimensions are significantly larger than $k$, ensuring that the condition specified in theorem 1 — that $N>k$ — is always met, making this theorem especially relevant for low-rank decompositions in feed-forward layers. For example the most common initialization schemes such as Kaiming (He et al., 2015) and Xavier (Glorot & Bengio, 2010) satisfy the requirements of our theorem.
@@
Figure 2: These figures display weight magnitudes for matrices with dimension $128\times 128$. The first figure shows a heatmap of a full-rank matrix initialized by Kaiming uniform, highlighting linear independence among rows. The second shows a low-rank matrix $\mathbf{W_{\text{lr}}}\,{=}\,\mathbf{UV}^{T}\,{\in}\,\mathbb{R}^{128{\times}128}$, with $\mathbf{U},\mathbf{V}\,{\in}\,\mathbb{R}^{128{\times}1}$ illustrating minimal linear independence. The final pair of figures reveal how applying a sine function element-wise, $\sin(\omega\cdot\mathbf{UV}^{T})$, with varying $\omega$, affects linear independence in low-rank matrices; specifically, $\omega=100$ and $\omega=2000$ progressively increase linear independence.
In figure 2 we give a visualization of our method in action. We consider a full-rank matrix, a low-rank matrix, and two sine-activated low-rank matrices with different frequencies. By visualizing the weight magnitudes in each matrix via a heatmap, we can clearly see how the sine-activated low-rank matrix increases rank and furthermore how increasing the frequency of the sine function increases the rank in accord with theorem 1.
@@ 4.1 Large Language Model
Setting. In the Transformer architecture, there are four weight matrices in the self-attention module ($\mathbf{W_{\text{q}}},\mathbf{W_{\text{k}}},\mathbf{W_{\text{v}}},\mathbf{W_{\text{o}}}$) and two in the MLP module($\mathbf{W_{\text{up}}},\mathbf{W_{\text{down}}}$). To evaluate RoBERTA V3, we follow up the LoRA architecture and implement low-rank adaptation only on $\mathbf{W_{\text{q}}}$ and $\mathbf{W_{\text{v}}}$. We study the performance of LoRA and sine LoRA in terms of different rank $k={1,2,4,8}$. To evaluate LLaMA 3-8B, we implement low-rank adaptations on five modules ($\mathbf{W_{\text{q}}},\mathbf{W_{\text{k}}},\mathbf{W_{\text{v}}},\mathbf{W_{\text{up}}},\mathbf{W_{\text{down}}}$) with different rank $k={4,8,16,32}$. For further implementation details see Section A.2.1.
@@
Analysis. Within the LoRA framework, featuring a low-rank multiplication component $\Delta\mathbf{W}=\mathbf{UV^{T}}$, we enhance this low-rank component with a sine function and assess the efficacy of our method. This adaptation amplifies the update significance due to the ‘intrinsic rank’ increase, facilitated by the sine-activation. Consequently, our approach attains superior performance at reduced rank levels $k$, compared to LoRA, effectively decreasing the count of learnable parameters.
@@
For comparisons to the recent DORA paper (Liu et al., 2024) see Section A.2.1.
@@ 4.2 Pretraining Vision Transformers
Figure 4: Low Rank ViT classification performance. Use of the sine-activation improves performance of the low-rank models, and even enables improvement relative to the Full Rank model.
@@
Analysis. Large models, such as ViT-Base, with an excessively large number of parameters, are prone to overfitting, where they perform well on training data but poorly on unseen data, especially when trained on relatively ‘small’ datasets like ImageNet-1k (Xu et al., 2024). Low-rank learning techniques can help in designing models that generalize better to new data by encouraging the model to learn more compact and generalizable representations to reduce overfitting. Additionally, while ViT architectures often underperform on smaller datasets, this method introduces a novel approach for efficiently training ViT models using small data collections. For frequency ablations in the rank 1 case, see Section A.2.2.
@@ 4.3 NeRF
Figure 5: (a) Using a non-transformed Low-Rank model leads to a complete loss of signal at extreme (rank $k=1$). In contrast, applying a sine-activation function is able to reconstruct details even at 1.3% of the Full-Rank parameters. (b) The Sine Low-Rank NeRF models show significant improvements across the rate-distortion curve relative to the Low-Rank models.
Analysis. NeRF models fit entire 3D scenes, and a high training PSNR leads to a high testing PSNR (Mildenhall et al., 2020). Employing structured weight matrices could result in a drop in performance due to the inherent constraints imposed by their structural design. Increasing the matrices’ rank enhances their memorization abilities significantly, especially when using a very low $k$. Starting from a low frequency, there is a rapid and consistent increase in PSNR. Consequently, as we elevate the rank level $k$, our results gradually align with the baseline NeRFs, which serve as the upper bound. For frequency ablations in the case of rank 1 and rank 5 see Section A.2.4.
@@
For this experiment, we evaluate performance on binary occupancy field reconstruction, which involves determining whether a given coordinate is occupied (Mescheder et al., 2019). Following (Saragadam et al., 2023), we sampled over a 512 $\times$ 512 $\times$ 512 grid with each voxel within the volume assigned a 1, and voxels outside the volume assigned a 0. We use the Thai Statue, Dragon and Lucy instance from the Stanford Scanning Repository. [^1] Available at https://graphics.stanford.edu/data/3Dscanrep/ We evaluate intersection over union (IoU) for the occupancy volumes. We used a coordinate-based MLP that includes two hidden layers, each with a width of 256 neurons, and employed the Gaussian activation function. The full-rank model achieves an accuracy of 97 (IoU). Figure 6 shows the 3D mesh representation of the Thai Statue, visualized using the low-rank method and the sine low-rank method for $k=1,2,5$. Applying the sine function to the low-rank matrix resulted in a significant enhancement and more precise shape delineation. Results on Dragon and Lucy are given in Table 13 in Section A.2.5. A frequency ablation in the rank 1 case is given in Section A.2.5.